# Tunneling dynamics of ${}^{164}$y supersolids and droplets

### v1

@@ Appendix B Linear ramping of the energy offset
Figure 10: Temporal evolution of the inter-well population imbalance, $\Delta\mathcal{N}(t)$, following a linear decrease of the energy offset from $D=7\hbar\omega_{x}/l_{{\rm osc}}$ to $D=0$ for different ramp-rates $\tau$ (see legend). A larger ramp-rate of the quench protocol leads progressively to tunneling suppression. The quasi-1D dBEC consists of $N=40000$ ${}^{164}$y atoms with $\epsilon_{{\rm dd}}=1.394$.
@@ Appendix C Numerical Scheme
To identify the correct ground state, we evaluate the corresponding energies of different initial states with precision up to ten decimal digits. The normalization of the wave function at every step of the imaginary time evolution is preserved by applying $\psi(\mathbf{r}^{\prime},t)\rightarrow N^{1/2}/{\norm{\psi(\mathbf{r}^{\prime},t)}}$. Convergence is justified by testifying that modifications of the wave function (at every grid point) between consecutive time-steps is lower than $10^{-6}$ and the corresponding energy alterations lie below $10^{-8}$. This serves as the initial state for the quench-induced tunneling dynamics which is monitored through real time propagation of the eGPE. The divergent behavior of the dipolar interaction potential at short distances is circumvented by transforming to momentum space Góral and Santos (2002) for the calculations. Subsequently, we employ the inverse Fourier transform to find the real space configurations relying on the convolution theorem. For the present simulations a 3D box with a spatial grid ($n_{x}\times n_{y}\times n_{z}$) is used which refers to ($1024\times 128\times 128$) in the quasi-1D case and ($512\times 512\times 128$) for the quasi-2D geometry. The spatial discretization is $\delta x=\delta y=\delta z=0.05~{}l_{{\rm osc}}$, and the time-step of the numerical integration $\delta t=10^{-5}/\omega_{x}$.

### v2

@@ Appendix B Linear ramping of the energy offset
Figure 10: Temporal evolution of the inter-well population imbalance, $\Delta\mathcal{N}(t)$, following a linear decrease of the energy offset from $D=7\hbar\omega_{x}/l_{{\rm osc}}$ to $D=0$ for different ramp-rates $\tau$ (see legend). A larger ramp-rate of the quench protocol leads progressively to tunneling suppression. The quasi-1D dBEC consists of $N=40000$ 164Dy atoms with $\epsilon_{{\rm dd}}=1.394$.
@@ Appendix C Numerical Scheme
To identify the correct ground state, we evaluate the corresponding energies of different initial states with precision up to ten decimal digits. The normalization of the wave function at every step of the imaginary time evolution is preserved by applying $\psi(\mathbf{r}^{\prime},t)\rightarrow N^{1/2}/{\norm{\psi(\mathbf{r}^{\prime},t)}}$. Convergence is justified by testifying that modifications of the wave function (at every grid point) between consecutive time-steps is lower than $10^{-6}$ and the corresponding energy alterations lie below $10^{-8}$. This serves as the initial state for the quench-induced tunneling dynamics which is monitored through real time propagation of the eGPE. The divergent behavior of the dipolar interaction potential at short distances is circumvented by transforming to momentum space Góral and Santos (2002) for the calculations. Subsequently, we employ the inverse Fourier transform to find the real space configurations relying on the convolution theorem. For the present simulations a 3D box with a spatial grid ($n_{x}\times n_{y}\times n_{z}$) is used which refers to ($1024\times 128\times 128$) in the quasi-1D case and ($512\times 512\times 128$) for the quasi-2D geometry. The spatial discretization is $\delta x=\delta y=\delta z=0.05~{}l_{{\rm osc}}$, and the time-step of the numerical integration $\delta t=10^{-5}/\omega_{x}$. Hence, in order to accurately simulate the real time dynamics of the system we guarantee that $(\delta t)^{2}<\delta x\delta y$  Antoine et al. (2013) is satisfied. Indeed, such spatial and time discretization steps ensure that the total particle number and total energy remain conserved, numerically of the order of $10^{-6}$, during the evolution time.